# Elephant random walks on infinite Cayley trees

### v2

@@ 4. Proofs of our main results
This completes the proof. ∎
Proof of Theorem 2.4.
The idea is to apply the martingale central limit theorem to $M_{n}=\sum_{k=1}^{n}Y_{k}$. We shall first compute the quadratic variation of $M_{n}$. Note that
$=\mathbb{E}[(\Delta_{k}-\Delta_{k-1})^{2}\mid\mathcal{F}_{k-1}]-G_{k-1}^{2}$
$=1-G_{k-1}^{2}.$
Thus
Also, since
we may rewrite the above result as
@@
$\leq\bigg(\frac{1-\frac{2}{d}-\varepsilon}{\frac{2(pd-1)}{d-1}}\bigg)^{-m}\cdot\mathbb{E}|\Xi_{n}|^{m}.$
Using the estimates from Lemma 3.3, with a choice of $m$ yielding linear decay for $\mathbb{E}|\Xi_{n}|^{m}$, we obtain a linear decay estimate for $\mathbb{P}\big(\frac{1}{n}\sum_{k=1}^{n}Z_{k}\leq 0\big)$ from (4).
@@
$\leq\bigg(\frac{1}{6}-\frac{1}{6n}-\varepsilon\bigg)^{-m}\cdot\mathbb{E}|\Xi_{n}|^{m},$
which, by Lemma 3.3, decays at least as fast as $n^{-1}$ for any $m>2$. Using this with (4), we obtain a linear decay estimate for $\mathbb{P}\big(\frac{1}{n}\sum_{k=1}^{n}Z_{k}\leq 0\big)$. ∎
@@
where, as before, $M_{n}=\sum_{k=1}^{n}Y_{k}$. For any $m\geq 1$, we have
$\mathbb{E}\bigg|\frac{\Delta_{n}}{n}$ $-\frac{d-2}{d}\bigg|^{m}$
$\leq 3^{m-1}\bigg[\frac{\mathbb{E}|M_{n}|^{m}}{n^{m}}+\bigg(\frac{2}{d}\bigg)^{m}\cdot\mathbb{E}\bigg(\frac{1}{n}\sum_{k=0}^{n-1}\mathds{1}(\Delta_{k}>0)\bigg)^{m}+\bigg|\frac{2(1-pd)}{d-1}\bigg|^{m}\cdot\mathbb{E}|\Xi_{n}|^{m}\bigg]$
$\lesssim\frac{\mathbb{E}|M_{n}|^{m}}{n^{m}}+\frac{1}{n}\sum_{k=0}^{n-1}\mathbb{P}(\Delta_{k}=0)+\mathbb{E}|\Xi_{n}|^{m}.$ (23)
@@
$\frac{\mathbb{E}|M_{n}|^{m}}{n^{m}}\leq 3^{m-1}\bigg[\mathbb{E}\bigg|\frac{\Delta_{n}}{n}-\frac{d-2}{d}\bigg|^{m}+\bigg(\frac{2}{d}\bigg)^{m}\cdot\frac{1}{n}\sum_{k=0}^{n-1}\mathbb{P}(\Delta_{k}=0)+\bigg|\frac{2(1-pd)}{d-1}\bigg|^{m}\cdot\mathbb{E}|\Xi_{n}|^{m}\bigg],$
@@
$\mathbb{E}\bigg|\frac{\Delta_{n}}{n}-\frac{d-2}{d}\bigg|^{m}\geq\frac{1}{3^{m-1}}\cdot\frac{\mathbb{E}|M_{n}|^{m}}{n^{m}}-\bigg(\frac{2}{d}\bigg)^{m}\cdot\frac{1}{n}\sum_{k=0}^{n-1}\mathbb{P}(\Delta_{k}=0)-\bigg|\frac{2(1-pd)}{d-1}\bigg|^{m}\cdot\mathbb{E}|\Xi_{n}|^{m}.$ (24)
@@
$\frac{1}{n}\sum_{k=0}^{n-1}\mathbb{P}(\Delta_{k}=0)\lesssim\begin{cases}\frac{1}{n}&\text{if }p\in[0,1/2)\text{ and }(p,d)\neq(0,3),\\ \frac{\log n}{n}&\text{otherwise.}\end{cases}$
@@
But, as shown in (16), $\frac{\langle M\rangle_{n}}{n}\xrightarrow{\text{a.s.}}\frac{4(d-1)}{d^{2}}$. Hence

### v3

@@ 4. Proofs of our main results
$\leq\exp\big(-C_{p,d,\varepsilon}\cdot r_{n}\big),$
where in the last line we have used Lemma 3.3, and $C_{p,d,\varepsilon}=\frac{1-\frac{2}{d}-\varepsilon}{\frac{2(pd-1)}{d-1}\widetilde{C}_{p,d}\exp(1)}$. Choosing a fixed $\varepsilon\in(0,1-\frac{2}{d})$, it now follows from (4) that
$\mathbb{P}\bigg(\frac{1}{n}\sum_{k=1}^{n}Z_{k}\leq 0\bigg)\lesssim\exp(-C_{p,d}\cdot r_{n}),$
for some constant $C_{p,d}>0$.
@@
$\lesssim\exp\big(-C_{0,3,\varepsilon}\cdot r_{n}\big),$
where in the last line we have used Lemma 3.3, and $C_{0,3,\varepsilon}=\frac{\frac{1}{7}-\varepsilon}{\widetilde{C}_{0,3}\exp(1)}$. Fix an $\varepsilon\in(0,\frac{1}{7})$. As in Case II, it follows from (4) that
$\mathbb{P}\bigg(\frac{1}{n}\sum_{k=1}^{n}Z_{k}\leq 0\bigg)\lesssim\exp(-C_{0,3}\cdot r_{n}),$
for some constant $C_{0,3}>0$. ∎
@@
where, as before, $M_{n}=\sum_{k=1}^{n}Y_{k}$. For any $m\geq 1$, by Minkowski’s inequality, we have
$\bigg[\mathbb{E}\bigg|\frac{\Delta_{n}}{n}$ $-\frac{d-2}{d}\bigg|^{m}\bigg]^{\frac{1}{m}}$
$\leq\frac{\big[\mathbb{E}|M_{n}|^{m}\big]^{\frac{1}{m}}}{n}+\frac{2}{d}\cdot\bigg[\mathbb{E}\bigg(\frac{1}{n}\sum_{k=0}^{n-1}\mathds{1}(\Delta_{k}=0)\bigg)^{m}\bigg]^{\frac{1}{m}}+\bigg|\frac{2(1-pd)}{d-1}\bigg|\cdot\big[\mathbb{E}|\Xi_{n}|^{m}\big]^{\frac{1}{m}}$
$\lesssim\frac{\big[\mathbb{E}|M_{n}|^{m}\big]^{\frac{1}{m}}}{n}+\bigg[\frac{1}{n}\sum_{k=0}^{n-1}\mathbb{P}(\Delta_{k}=0)\bigg]^{\frac{1}{m}}+\big[\mathbb{E}|\Xi_{n}|^{m}\big]^{\frac{1}{m}}.$ (22)
@@
$\frac{\big[\mathbb{E}|M_{n}|^{m}\big]^{\frac{1}{m}}}{n}\leq\bigg[\mathbb{E}\bigg|\frac{\Delta_{n}}{n}-\frac{d-2}{d}\bigg|^{m}\bigg]^{\frac{1}{m}}+\frac{2}{d}\cdot\bigg[\frac{1}{n}\sum_{k=0}^{n-1}\mathbb{P}(\Delta_{k}=0)\bigg]^{\frac{1}{m}}+\bigg|\frac{2(1-pd)}{d-1}\bigg|\cdot\big[\mathbb{E}|\Xi_{n}|^{m}\big]^{\frac{1}{m}},$
@@
$\bigg[\mathbb{E}\bigg|\frac{\Delta_{n}}{n}-\frac{d-2}{d}\bigg|^{m}\bigg]^{\frac{1}{m}}\geq\frac{\big[\mathbb{E}|M_{n}|^{m}\big]^{\frac{1}{m}}}{n}-\frac{2}{d}\cdot\bigg[\frac{1}{n}\sum_{k=0}^{n-1}\mathbb{P}(\Delta_{k}=0)\bigg]^{\frac{1}{m}}-\bigg|\frac{2(1-pd)}{d-1}\bigg|\cdot\big[\mathbb{E}|\Xi_{n}|^{m}\big]^{\frac{1}{m}}.$ (23)
@@
$\frac{1}{n}\sum_{k=0}^{n-1}\mathbb{P}(\Delta_{k}=0)\lesssim\frac{1}{n}.$
@@
But, as shown in (25) (in the proof of Proposition 2.1), $\frac{\langle M\rangle_{n}}{n}\xrightarrow{\text{a.s.}}\frac{4(d-1)}{d^{2}}$. Hence
@@
Proof of Proposition 2.1.
The proof is an application of the martingale central limit theorem to $M_{n}=\sum_{k=1}^{n}Y_{k}$. We shall first compute the quadratic variation of $M_{n}$. Note that
$=\mathbb{E}[(\Delta_{k}-\Delta_{k-1})^{2}\mid\mathcal{F}_{k-1}]-G_{k-1}^{2}$
$=1-G_{k-1}^{2}.$
Thus
Also, since
we may rewrite the above result as
This completes the proof. ∎
5. Conclusion
In conclusion, we have introduced a generalisation of the elephant random walk of [ST04] to finite generated groups and their Cayley graphs. Focusing on the Bethe lattice, we have shown, by leveraging the connection of the ERW dynamics with urn models first observed in [BB16], that memory has no effect on the first order behaviour of the asymptotic speed, thus matching speed of the simple random walk. On the other hand, the rate of convergence to the limiting speed appears to undergo a phase transition, as evidenced by our upper bounds and numerical experiments. Obtaining precise limit distributions is an interesting open question, as is proving exponential decay of the return probability in all memory regimes. Answering these questions would potentially require a more careful decoupling of the underlying urn process governing the step counts from the geometry of the Cayley graph. Investigation of how memory interacts with geometry in other finitely generated infinite groups is another intriguing question, which we hope to return to in the future.